(8) The country described in the table below produces and consumes only three goods: ice cream, chocolate cake, and t-shirts. The quantity produced and price of each good in 2016 and 2017 are provided.Price 2016 Quantity 150 600 2017 Quantity 170 Price $22 $1.50 $20 $2 Good t-shirts Ice cream (scoops Chocolate cake (slices)
Answer #1

Nominal GDP = current year quantity*current year price

Real GDP = current year quantity*base year price

a) Nominal GDP in 2016 = 150*22+600*1.5+1200*2.25 = 3300+900+2700 = 6900

b) Nominal GDP in 2017 = 170*20+600*2+1100*3 = 3400+1200+3300 = 7900

c) 2017 as the base year, the Real GDP in 2016 = 150*20+600*2+1200*3 = 3000+1200+3600 = 7800

d) 2017 as the base year, the Real GDP in 2017 = 7900

Percentage growth rate = 7900-7800/7800*100 = 1.28%

As per HOMEWORKLIB RULES, the first question is answered